# Simulating Online Social Media Conversations on Controversial Topics Using AI Agents Calibrated on Real-World Data

### v1

@@ Abstract
Online social networks offer a valuable lens to analyze both individual and collective phenomena. Researchers often use simulators to explore controlled scenarios, and the integration of Large Language Models (LLMs) makes these simulations more realistic by enabling agents to understand and generate natural language content. In this work, we investigate the behavior of LLM-based agents in a simulated microblogging social network. We initialize agents with realistic profiles calibrated on real-world online conversations from the 2022 Italian political election and extend an existing simulator by introducing mechanisms for opinion modeling. We examine how LLM agents simulate online conversations, interact with others, and evolve their opinions under different scenarios. Our results show that LLM agents generate coherent content, form connections, and build a realistic social network structure. However, their generated content displays less heterogeneity in tone and toxicity compared to real data. We also find that LLM-based opinion dynamics evolve over time in ways similar to traditional mathematical models. Varying parameter configurations produces no significant changes, indicating that simulations require more careful cognitive modeling at initialization to replicate human behavior more faithfully. Overall, we demonstrate the potential of LLMs for simulating user behavior in social environments, while also identifying key challenges in capturing heterogeneity and complex dynamics.
1 Introduction
@@
Online social networks have evolved beyond mere communication platforms, becoming digital arenas where users express emotions, form opinions, and shape behaviors (Bakshy, Messing, and Adamic 2015). These environments provide unique opportunities to study complex collective phenomena such as polarization, content diffusion, and large-scale social dynamics (Vosoughi, Roy, and Aral 2018). The widespread availability of digital traces, enabled by the pervasive use of online technologies, has fueled the rise of computational social science (Lazer et al. 2009), which seeks to explain human behavior and social processes through computational methods.
@@
A prominent approach in this field involves the use of simulation tools (Squazzoni, Jager, and Edmonds 2014). Simulations make it possible to create controlled virtual environments where researchers can test hypotheses, compare strategies, and observe the evolution of user behavior under conditions that would be difficult—or ethically problematic—to reproduce in the real world (Rossetti et al. 2024). For example, one can investigate the spread of harmful content and rumors (Hu et al. 2025) or assess how different recommendation algorithms shape user activity (Törnberg et al. 2023), all without intervening directly on live platforms. Despite their promise, building realistic simulations of online social networks remains challenging. Emergent behaviors in these systems are driven by numerous individual-level factors that are difficult to predict or formalize. Human interactions involve ambiguity, context-dependence, and variability, which complicate the design of models capable of capturing social complexity (Gao et al. 2023). Agent-Based Modeling (ABM) has long been employed to address this challenge. ABMs represent systems as collections of autonomous agents, each following a set of predefined behavioral rules (Macy and Willer 2002). This bottom-up approach is particularly well suited for modeling online platforms, where heterogeneous individuals interact locally and generate emergent global patterns (Gausen, Luk, and Guo 2021). ABMs thus allow researchers to explore the interplay between micro-level interactions and macro-level outcomes. Traditional ABMs, however, typically rely on simplified behavioral rules (Conte and Paolucci 2014). While such models have yielded important insights, they struggle to capture the richness of human behavior, which is mediated by language, emotions, and social context (Törnberg et al. 2023).
@@
In this context, Large Language Models (LLMs) represent a promising extension of ABM. Unlike traditional agents, which follow fixed and narrow rules, LLM-based agents can generate nuanced, coherent, and context-aware behaviors (Park et al. 2024; Fontana, Pierri, and Aiello 2025). Their capacity to simulate conversations, express emotions, adopt perspectives, and deploy diverse interaction strategies enables them to approximate human-like behavior with unprecedented fidelity (Park et al. 2023; Corso et al. 2025). Moreover, LLM agents can be enriched with persistent traits—such as personality, ideology, or memory of past interactions—that allow them to act consistently across time (Rossetti et al. 2024). This makes them particularly powerful for reproducing both individual-level realism and emergent collective phenomena (Park et al. 2023). Early studies have demonstrated the potential of LLM-driven simulations, suggesting that these models offer a promising avenue for advancing the study of online behavior and warrant systematic exploration (Gao et al. 2023; Törnberg et al. 2023; Rossetti et al. 2024).
@@
This work builds on Y Social (Rossetti et al. 2024), a social media simulator that reproduces online platforms within a controlled environment. In this simulator, LLM-powered agents behave like users: they consume content, post, and interact with one another. Using a data-driven approach that leverages Y Social framework with real-world online conversations from the 2022 Italian election we aim to address the following research questions:
@@
To this end, our contribution extends Y in two key directions. First, we seed agents with activity patterns and political leaning drawn from a dataset of real-world Twitter conversations collected around the 2022 Italian elections (Pierri, Liu, and Ceri 2023), enhancing the realism of their behavior and enabling systematic comparison with an empirically grounded scenario. Second, we equip agents with opinions on specific political topics, and observe the dynamics of opinion change over time through interactions. We conduct a series of experiments varying key parameters such as the LLM powering agents, the network structure and the recommendation system. We then compare simulated outcomes against real-world data to assess both the validity and the limitations of LLM-based social simulations. By doing so, this study both evaluates the potential of LLMs as social agents and identifies the challenges that must be addressed to use them as reliable tools for modeling social behavior.
2 Related work
2.1 Simulating social networks
Social simulations have been widely used to study group behavior and the nonlinear effects of individual interactions (Squazzoni, Jager, and Edmonds 2014). Agent-Based Modeling (ABM) focuses on local agent dynamics and demonstrates how simple interactions can reproduce complex social phenomena (Macy and Willer 2002). ABMs also allow the linking of micro- and macro-level phenomena, highlighting causal relationships between individual behavior and network structure (Squazzoni, Jager, and Edmonds 2014). A key limitation of traditional ABMs is the simplicity of agent behavior rules (Conte and Paolucci 2014) and their limited capacity for realistic social interaction (Törnberg et al. 2023). Recent advances in AI and Large Language Models (LLMs) offer a way to overcome these constraints, enabling agents to engage in realistic conversations and exhibit human-like behaviors (Park et al. 2023). Several recent studies have explored the use of LLM agents in simulated social network environments. Below, we discuss three representative simulators.
Törnberg et al. (2023) simulated three social media platforms, each with a distinct content recommendation algorithm, to assess how news feed personalization affects conversation quality and cross-party interactions. Agents, powered by LLMs, were initialized with demographic characteristics, political leanings, interests, and attitudes from the 2020 American National Election Study (ANES). The first platform promoted popular posts from followed users, while the second suggested globally popular posts; both reduced cross-party interactions and increased toxicity. In contrast, a third “bridging” algorithm recommended posts popular among users with opposing views, resulting in more constructive, less toxic, and more inter-partisan interactions. This work underscores the significant influence of recommender systems on online discourse quality.
@@ 2.1 Simulating social networks
Rossetti et al. (2024) introduced Y, a social media digital twin, a system designed to digitally replicate a real-world system to allow analysis, simulation and experimentation in a controlled environment. The users of these simulations are LLM agents, and they can perform all the common actions available on the most popular social media, including posting, commenting, replying, reacting and following other users. Other modules also allow the integration of images. User profiles are enriched with attributes including their interests, political leaning, demographic data and personality, which is defined according to the Big Five model (Barrick and Mount 1991; McCrae and John 1992). To make the simulations even more realistic, Y also includes the possibility of adding external input to the simulation. Specifically, users can share news gathered from selected websites, provided through RSS (Really Simple Syndication) feeds. Moreover, Y includes the implementation of various recommender and ranking algorithms to promote specific content or users. This enables further study of the impact the algorithmic curation has on online conversations and users’ behavior. This expands the approach of Törnberg et al. (2023), by offering a more flexible and realistic simulation framework.
2.2 Opinion Dynamics
@@
The Friedkin–Johnsen model (Friedkin and Johnsen 1990) addresses this by introducing a susceptibility parameter, allowing agents to retain part of their initial opinion. More recent extensions incorporate state-dependent updating, where adjustments depend on current beliefs rather than the initial stance (Ye, Liu, and Anderson 2018; Liu et al. 2018).
While mathematical models provide valuable abstractions, they reduce opinions to numerical values and overlook elements such as language, tone, and personality. Recent studies address this by employing LLMs as agents, which can impersonate profiles, engage in realistic interactions, and express beliefs in natural language. For example, Cau et al. (2025) simulated paired discussions on the Ship of Theseus paradox, a topic without factual resolution. Agents held discrete opinions (0–6) and updated them stepwise if persuaded. Results showed that LLMs often aligned with partners’ views, but the setup underutilized LLM capabilities, as agents lacked richer personalization such as demographics or personality traits.
@@ 3 Experimental Design
  • Demographics & personality. Agents are assigned the following attributes: age (integer, sampled from weighted 2024 X statistics, range 18–60), gender (categorical; sampled to match platform-weighted proportions), and nationality (fixed to Italian to match the case study). These attributes are used as input features to the prompts and affect probabilities for actions such as posting, commenting, or following, but also unfollowing and, more in general, all the possible interactions that the agents can perform, together with the internal opinion-updating mechanism. We leave more details in Appendix B,B.2.
  • Political leaning & coalition principles. Each agent receives a coalition label (i.e., Right, Centre-Left, Third Pole, M5S) sampled from the source dataset and inferred based on the retweeting behaviour with respect to representatives of different parties Pierri, Liu, and Ceri (2023). For each coalition, we store a short, standardized principles string summarizing its typical policy priorities and rhetorical framing (used as contextual prompt material for the LLM). Political leaning is represented categorically (coalition), whereas topic opinions are represented on a continuous range to seed initial opinion values. More on this in Appendix C.
@@
To make our agents more realistic, some attributes have been initialized based on the ITA-ELECTION-22 dataset (Pierri, Liu, and Ceri 2023), a collection of Twitter posts in Italian language around the Italian political election in 2022. Specifically, the attributes initialized from the dataset in this work are: the political leaning, the average toxicity of posts and comments posted on the platform (estimated using Detoxify), and the activity level, for each user. The activity is computed by converting the number of tweets posted by each user into a continuous value in the range $[0,1]$, with a logaritmic normalization to reduce the impact of outliers. The formula used is the following:
@@
The agents create the network connections over time via the follow interaction, allowing the network structure to emerge dynamically as the simulation progresses. In the extended framework, the opinion update is directly performed by LLMs. The evolving network structure was then used to compute the mathematical equivalent of agents’ stances through the Friedkin–Johnsen model, which we executed in parallel to our simulation, in order to compare the results of these two approaches.
At the beginning of the simulation, a population of agents is initially generated. At this stage the agents may be already be connected with each other, depending on the network initialization strategy. However, throughout the simulation, agents have the possibility to create new links or remove the existing ones, evaluating the interactions they had with other users. In this way, the network structure dynamically evolves over time according to the agents’ behavior and interactions. In Procedure 1 we then describe each step of a day in the simulation.
@@
At the end of the day, active agents are asked to update their opinion on the topics they discussed. This phase is critical to study the opinion dynamics: it makes it possible to observe how social interactions and the received content impact the evolution of individual views. We normalized all numeric attributes are (facilitating their use in probabilistic decision functions), and we stored textual fields (principles, justifications, topic descriptions) as compact prompt templates to ensure consistent LLM conditioning. The behavioral Decision(profile, content) function combines numeric traits, coalition priors, and content-topic alignment to produce an action probability distribution.
@@
The discussions between agents focused on four major topics of debate that characterized the Italian 2022 election, selected for their political salience and the absence of a single “ground truth”. These topics were selected because they represent politically relevant issues in the Italian context of 2022, on which the main coalitions held different positions. This allows the simulation to generate meaningful political discussions and potential conflicts among agents. Furthermore, these issues don’t have a single ground truth, meaning that the simulations do not necessarily lead to consensus (Cau et al. 2025).

### v2

@@ Abstract
Online social networks offer a valuable lens to analyze both individual and collective phenomena. Researchers often use simulators to explore controlled scenarios, and the integration of Large Language Models (LLMs) makes these simulations more realistic by enabling agents to understand and generate natural language content. In this work, we offer an exploratory investigation of the behavior of LLM-based agents in a simulated microblogging social network. We initialize agents with realistic profiles and opinions calibrated on real-world online conversations from the 2022 Italian political election and extend an existing simulator by introducing mechanisms for opinion modeling. We examine how LLM agents simulate online conversations, interact with others, and evolve their opinions under different scenarios. Our results show that LLM agents generate coherent content, form connections, and build a realistic social network structure. However, their generated content displays less heterogeneity in tone and toxicity compared to real data. We also find that LLM-based opinion dynamics evolve over time in ways similar to traditional mathematical models. Varying parameter configurations produces no significant changes, indicating that simulations require more careful cognitive modeling at initialization to replicate human behavior more faithfully. Overall, we demonstrate the potential of LLMs for simulating user behavior in social environments, while also identifying key challenges in capturing heterogeneity and complex dynamics.
Introduction
@@
Online social networks have evolved beyond mere communication platforms, becoming digital arenas where users express emotions, form opinions, and shape behaviors (Bakshy et al. 2015). These environments provide unique opportunities to study complex collective phenomena such as polarization, content diffusion, and large-scale social dynamics (Vosoughi et al. 2018). The widespread availability of digital traces, enabled by the pervasive use of online technologies, has fueled the rise of computational social science (Lazer et al. 2009), which seeks to explain human behavior and social processes through computational methods.
@@
A prominent approach in this field involves the use of simulation tools (Squazzoni et al. 2014). Simulations make it possible to create controlled virtual environments where researchers can test hypotheses, compare strategies, and observe the evolution of user behavior under conditions that would be difficult—or ethically problematic—to reproduce in the real world (Rossetti et al. 2024). For example, one can investigate the spread of harmful content and rumors (Hu et al. 2025) or assess how different recommendation algorithms shape user activity (törnberg2023evaluate), all without intervening directly on live platforms. Despite their promise, building realistic simulations of online social networks remains challenging. Emergent behaviors in these systems are driven by numerous individual-level factors that are difficult to predict or formalize. Human interactions involve ambiguity, context-dependence, and variability, which complicate the design of models capable of capturing social complexity (Gao et al. 2023). Agent-Based Modeling (ABM) has long been employed to address this challenge. ABMs represent systems as collections of autonomous agents, each following a set of predefined and simplified behavioral rules (Macy and Willer 2002; Conte and Paolucci 2014). While such models have yielded important insights, they struggle to capture the richness of human behavior, which is mediated by language, emotions, and social context (törnberg2023evaluate).
@@
In this context, Large Language Models (LLMs) represent a promising extension of ABM. Unlike traditional agents, which follow fixed and narrow rules, LLM-based agents can generate nuanced, coherent, and context-aware behaviors (Park et al. 2024; Fontana et al. 2025). Their capacity to simulate conversations, express emotions, adopt perspectives, and deploy diverse interaction strategies enables them to approximate human-like behavior with unprecedented fidelity (Park et al. 2023; Corso et al. 2025a). Moreover, LLM agents can be enriched with persistent traits—such as personality, ideology, or memory of past interactions—that allow them to act consistently across time (Rossetti et al. 2024). This makes them particularly powerful for reproducing both individual-level realism and emergent collective phenomena (Park et al. 2023). Early studies have demonstrated the potential of LLM-driven simulations, suggesting that these models offer a promising avenue for advancing the study of online behavior and warrant systematic exploration (Gao et al. 2023; törnberg2023evaluate; Rossetti et al. 2024).
@@
This work builds on Y Social (Rossetti et al. 2024), a social media simulator that reproduces online platforms within a controlled environment. In this simulator, LLM-powered agents behave like users: they consume content, post, and interact with one another. Using a data-driven approach that leverages Y Social framework with real-world online conversations from the 2022 Italian election, we aim to address the following research questions:
@@
To this end, our contribution extends Y in two key directions. First, we seed agents with activity patterns and political leaning drawn from a dataset of real-world Twitter conversations collected around the 2022 Italian elections (Pierri et al. 2023), enhancing the realism of their behavior and enabling systematic comparison with an empirically grounded scenario. Second, we equip agents with opinions on specific political topics and observe the dynamics of opinion change over time through interactions. We conduct a series of experiments varying key parameters such as the LLM powering agents, the network structure, and the recommendation system. We then compare simulated opinion dynamics outcomes against real-world data to assess both the validity and the limitations of LLM-based social simulations. By doing so, this study both explores the potential of LLMs as social agents and identifies the challenges that must be addressed to use them as reliable tools for modeling social behavior.
Related Work
Simulating social networks
Social simulations have been widely used to study group behavior and the nonlinear effects of individual interactions (Squazzoni et al. 2014). Agent-Based Modeling (ABM) focuses on local agent dynamics and demonstrates how simple interactions can reproduce complex social phenomena (Macy and Willer 2002). ABMs also allow the linking of micro- and macro-level phenomena, highlighting causal relationships between individual behavior and network structure (Squazzoni et al. 2014). A key limitation of traditional ABMs is the simplicity of agent behavior rules (Conte and Paolucci 2014) and their limited capacity for realistic social interaction (törnberg2023evaluate). Recent advances in AI and Large Language Models (LLMs) offer a way to overcome these constraints, enabling agents to engage in realistic conversations and exhibit human-like behaviors (Park et al. 2023). Several recent studies have explored the use of LLM agents in simulated social network environments. Below, we discuss three representative simulators.
törnberg2023evaluate simulated three social media platforms, each with a distinct content recommendation algorithm, to assess how news feed personalization affects conversation quality and cross-party interactions. Agents, powered by LLMs, were initialized with demographic characteristics, political leanings, interests, and attitudes from the 2020 American National Election Study (ANES). The first platform promoted popular posts from followed users, while the second suggested globally popular posts; both reduced cross-party interactions and increased toxicity. In contrast, a third “bridging” algorithm recommended posts popular among users with opposing views, resulting in more constructive, less toxic, and more inter-partisan interactions. This work underscores the significant influence of recommender systems on online discourse quality.
@@
Rossetti et al. (2024) introduced Y, a social media digital twin, a system designed to digitally replicate a real-world system to allow analysis, simulation, and experimentation in a controlled environment. The users of these simulations are LLM agents, and they can perform all the common actions available on the most popular social media, including posting, commenting, replying, reacting, and following other users. Other modules also allow the integration of images. User profiles are enriched with attributes including their interests, political leaning, demographic data, and personality, which is defined according to the Big Five model (Barrick and Mount 1991; McCrae and John 1992). To make the simulations even more realistic, Y also includes the possibility of adding external input to the simulation. Specifically, users can share news gathered from selected websites, provided through RSS (Really Simple Syndication) feeds. Moreover, Y includes the implementation of various recommender and ranking algorithms to promote specific content or users. This enables further study of the impact the algorithmic curation has on online conversations and users’ behavior. This expands the approach of törnberg2023evaluateby offering a more flexible and realistic simulation framework.
Opinion Dynamics
@@
The Friedkin–Johnsen model (Friedkin and Johnsen 1990) addresses this by introducing a susceptibility parameter, allowing agents to retain part of their initial opinion. More recent extensions incorporate state-dependent updating, where adjustments depend on current beliefs rather than the initial stance (Ye et al. 2018; Liu et al. 2018).
While mathematical models provide valuable abstractions, they reduce opinions to numerical values and overlook elements such as language, tone, and personality. Recent studies address this by employing LLMs as agents, which can impersonate profiles, engage in realistic interactions, and express beliefs in natural language. For example, Cau et al. (2025) simulated paired discussions on the Ship of Theseus paradox, a topic without factual resolution. Agents held discrete opinions (0–6) and updated them stepwise if persuaded. Results showed that LLMs often aligned with partners’ views, but the setup underutilized LLM capabilities, as agents lacked richer personalization, such as demographics or personality traits.
@@
  • Demographics. Agents are assigned the following attributes: age (integer, sampled from weighted 2024 X statistics, range 18–60), gender (categorical; binary variable, sampled with probabilities proportional to the gender distribution observed on the real platform based on Statista (Statista Research Department 2024)), and nationality (fixed to Italian to match the case study). These attributes are used as input features to the prompts and affect probabilities for actions such as posting, commenting, or following, but also unfollowing and, more in general, all the possible interactions that the agents can perform, together with the internal opinion-updating mechanism. We leave more details in Appendix A.
  • Political leaning & coalition principles. Each agent receives a coalition label (i.e., Right, Centre-Left, Third Pole, M5S) sampled from the source dataset and inferred based on the retweeting behaviour with respect to representatives of different parties (Pierri et al. 2023). For each coalition, we store a short, standardized principles string summarizing its typical policy priorities and rhetorical framing (used as contextual prompt material for the LLM). Political leaning is represented categorically (coalition), whereas topic opinions are represented on a continuous range to seed initial opinion values. Additional information on the coalitions can be found in Appendix A.
@@
To make our agents more realistic, some attributes have been initialized based on the ITA-ELECTION-22 dataset (Pierri et al. 2023), a collection of Twitter posts in the Italian language around the Italian political election in 2022. Specifically, the attributes initialized from the dataset in this work are: the political leaning, the average toxicity of posts and comments posted on the platform (estimated using Detoxify), and the activity level, for each user. The activity is computed by converting the number of tweets posted by each user into a continuous value in the range $[0,1]$, with a logarithmic normalization to reduce the impact of outliers. The formula used is the following:
@@
These two initialization strategies represent the two extreme regimes of social connectivity: the Empty Network captures a scenario with no social influence, where users’ behavior is driven solely by individual cognition and content exposure, while the Fully Connected Network represents the maximal influence regime, where every user can directly observe and react to every other user. Together, these extremes represent the upper and lower bound the spectrum of possible interaction structures, allowing us to isolate the effect of social connectivity itself on emergent dynamics. The agents create the network connections over time via the follow interaction, allowing the network structure to emerge dynamically as the simulation progresses. In the extended framework, the opinion update is directly performed by LLMs. The evolving network structure was then used to compute the mathematical equivalent of agents’ stances through the Friedkin–Johnsen (FJ) model, which we executed in parallel to our simulation, in order to compare the results of these two approaches. We chose to rely on the FJ model as our baseline since it is widely used in Agent-based modeling literature (Sun and Zhang 2023; Disarò and Valcher 2024) and, even in its original version, is easily comparable to our LLM-based setup.
At the beginning of the simulation, a population of agents is initially generated. At this stage, the agents may already be connected with each other, depending on the network initialization strategy. However, throughout the simulation, agents have the possibility to create new links or remove the existing ones, evaluating the interactions they had with other users. In this way, the network structure dynamically evolves over time according to the agents’ behavior and interactions. In Procedure 1, we then describe each step of a day in the simulation.
@@
At the end of the day, active agents are asked to update their opinion on the topics they discussed. This phase is critical to study the opinion dynamics: it makes it possible to observe how social interactions and the received content impact the evolution of individual views. We normalized all numeric attributes (facilitating their use in probabilistic decision functions), and we stored textual fields (principles, justifications, topic descriptions) as compact prompt templates to ensure consistent LLM conditioning. The behavioral Decision(profile, content) function combines numeric traits, coalition priors, and content-topic alignment to produce an action probability distribution.
@@
The discussions between agents focused on four major topics of debate that characterized the Italian 2022 election, selected for their political salience and the presence of a range of stances for each of these issues. These topics were selected because they represent politically relevant issues in the Italian context of 2022, on which the main coalitions held different positions. This allows the simulation to generate meaningful political discussions and potential conflicts among agents. Furthermore, since these issues are characterized by many different stances, the simulations do not necessarily lead to consensus (Cau et al. 2025).